A sample of 100 apparently normal adult males, 25 years old, had a mean systolic blood pressure of
125. It is believed that the population standard deviation is 15.
1
Expert's answer
2020-04-14T18:20:30-0400

"n=100,\bar{X}=25,\sigma=15."


"CI=(\bar{X}-z_{1-\alpha/2}\cdot{\sigma\over\sqrt{n}}, \bar{X}+z_{1-\alpha/2}\cdot{\sigma\over\sqrt{n}})"

(a) Construct a 97% confidence interval.

"\alpha=0.03, z_{1-\alpha/2}=2.17"


"CI=(125-2.17\cdot{15\over\sqrt{100}}, 125+2.17\cdot{15\over\sqrt{100}})="

"=(121.745, 128.255)"

Therefore, based on the data provided, the 97% confidence interval for the population mean is "121.745<\mu<128.255" , which indicates that we are 97%  confident that the true population "\mu"  is contained by the interval "(121.745, 128.255)."


(b) How many adult males must we select so that the margin of error is not greater than 2.5 kg

considering the confidence level of 98%?

"\alpha=0.02, z_{1-\alpha/2}=2.326"


"margin\ of\ error=z_{1-\alpha/2}\cdot{\sigma\over\sqrt{n}}\leq E=2.5"

"N\geq({z_{1-\alpha/2}\cdot\sigma \over E})^2"

"N\geq({2.326\cdot15\over 2.5})^2"

"N\geq195"

"195" adults.